# Very Massive Stars and High N/O: A Tale of the Nitrogen-enriched Super Star Cluster in the Sunburst Arc

### v1

@@ 3.4 Simulation suite
The star formation physics adopted in this study is FIRE-3 (Hopkins et al., 2023), which effectively evaluates the star formation rate of the GMC (while the underlying IMF is not resolved but modeled with the method in § 3.1). Each gas cell will turn into a FIRE SSP particle only when 1) its density is above a threshold ($1000\,\rm cm^{-3}$); 2) it is self-gravitating despite turbulence, thermal energy, and magnetic fields; 3) there is convergent inflow ($\nabla\cdot\mathbf{v}<0$); 4) its Jeans mass is smaller than the cell’s mass (Grudić et al., 2021). Moreover, the star formation rate is normalized with the molecular fraction of the cell (Hopkins et al., 2018a). Once such an SSP particle forms, Poisson sampling is performed to determine the number of VMSs, and individual VMS particles are spawned accordingly (for details see Shi & Murray, 2025).
@@
Radiative (in UV, Optical, IR) and kinetic (stellar winds and supernovae) feedback are treated following Hopkins et al. (2018b, a). Radiation transfer is treated in the same way as described in Hopkins et al. (2012) and Hopkins et al. (2018a), i.e., the luminosity in different radiation bands $L_{\nu}$ is first attenuated by gas and dust in the vicinity of the stars before it is transmitted to the ISM, while the attenuation also causes momentum coupling, radiative heating, and cooling to the nearby gas. In particular, in the UV band where massive stars are bright, the flux-weighted opacity is $\kappa_{\nu}=1800\,(Z/Z_{\odot})\,{\rm cm}^{2}\,{\rm g}^{-1}$ (i.e., dust abundance is proportional to metallicity). For VMS, luminosities in different bands are treated in the same way as those of single stars in Grudić et al. (2021) by assuming black-body stellar spectra.
The code deals with the mixing of metals by treating them as passive scalars and solving the advection-diffusion equation with source terms (i.e., yields from stars), as implemented in Hopkins et al. (2018a) and Rennehan et al. (2019) (and verified in Colbrook et al., 2017). Apart from the continuity equation that tracks the resolved mixing in the simulation, the transport of metals due to unresolved, small-scale turbulence is simulated using the Smagorinsky sub-grid turbulent diffusion model (Smagorinsky, 1963; Escala et al., 2018).
The initial conditions for our simulations are generated in the same way as several previous works (Grudić et al., 2018; Shi & Murray, 2025). The GMC is initially spherical, uniform in density, and non-spinning. The gas cells are initialized with a random turbulent velocity with 50% of the modes solenoidal, and the total turbulent energy equals the gravitational binding energy, i.e., $\alpha_{\rm vir}\equiv 2\,T/|U|=2$. While recent observations of molecular clouds (e.g., Sun et al., 2020) found significant scattering in $\alpha_{\rm vir}$ distribution, $\alpha_{\rm vir}=2$ is a typical value for massive GMCs (Chevance et al., 2023). The initial magnetic field is uniform, and the magnetic energy equals 1% of the gravitational binding energy. The spatially uniform density and magnetic field will become turbulent soon after the simulation starts due to the turbulent velocity field. Elemental abundances of the GMC are initialized following the empirical fitting relations Eqs. (3) and (4).
@@
Each of our simulation runs are labeled with M%g_R%g_Z%g, encoding the GMC’s initial mass $M_{\rm cloud}$ (in $M_{\odot}$), radius $R_{\rm cloud}$ (in pc), and metallicity $Z$ (in $Z_{\odot}$), respectively (e.g., M3e7_R70_Z0.2 means $M_{\rm cloud}=3\times 10^{7}\,M_{\odot}$, $R_{\rm cloud}=70\,{\rm pc}$, and $Z=0.2\,Z_{\odot}$). All GMCs are initialized with $64^{3}$ equal-mass gas cells and evolve for a physical duration of 10 Myr. When VMSs are modeled, a Kroupa IMF that extends to $m_{\rm max}=300\,M_{\odot}$ is assumed, and stars more massive than $m_{\rm cut}=75\,M_{\odot}$ are explicitly treated as individual stars, while less massive ones are collectively modeled as FIRE SSPs. We run four groups of simulations for different goals, as listed below (also see Table 1).
@@
    Fiducial simulations. We use the FIRE+VMS module explained earlier to perform the fiducial simulations. Based on properties of the LyC leaking cluster (Pascale et al., 2023), our fiducial simulations have fixed $M_{\rm cloud}=3\times 10^{7}\,M_{\odot}$ but vary $R_{\rm cloud}$ ($60,70,80\,\rm pc$) and $Z/Z_{\odot}$ (0.2, 0.15, 0.25). As will be explained in § 4, these simulations can (roughly) reproduce the observations.
@@
    GMC surface density tests. This variation keeps the fiducial setup but initializes with different $M_{\rm cloud}\,(10^{7},10^{8}\,\rm M_{\odot})$, while keeping the same $R_{\rm cloud}$ ($60,70,80\,\rm pc$) and $Z/Z_{\odot}=0.2$. These simulations explore the impact of the GMC initial mass (and surface density), to accommodate the order-of-unity uncertainty in the star cluster’s mass due to the magnification factor of gravitational lensing (Pascale et al., 2023).
@@
    FIRE-only simulations. These simulations remove explicitly-modeled VMSs and utilize FIRE-3 physics (where the underlying IMF only extends to $100\,M_{\odot}$) to highlight the effect of VMSs in shaping the star formation, especially the chemical evolution. They initialize with $M_{\rm cloud}=3\times 10^{7}\,M_{\odot}$, $R_{\rm cloud}=60,70,80\,\rm pc$, and $Z=0.2\,Z_{\odot}$. In particular, since we removed “failed SNe” from FIRE-3 when implementing the hybrid FIRE+VMS model (§ 3.3), we also run such a version of simulations in contrast to the standard FIRE-3 physics.
@@
4 Results
Figure 6: Star formation and the chemical abundance of high-pressure ($10^{9}\,\rm K\,cm^{-3}$) gas throughout the evolution of our “fiducial” group of simulations, as a function of “time since star formation.” These simulations initialize with different GMCs (M3e7_R60, M3e7_R70, M3e7_R80) and metallicities: $0.2\,Z_{\odot}$ (top row), $0.15\,Z_{\odot}$ (second row), and $0.25\,Z_{\odot}$ (bottom row). Each row, from left to right, shows the 1) star formation history, in general (solid), or of the most massive bound cluster (dotted), or of all VMSs (dot dashed); 2) UV half-light radius (of all stars in the field; solid) and the 3D half-mass radius (of the most massive bound cluster; dotted); 3) median $\log(\rm N/O)$ for the high-pressure gas; 4) median $\log(\rm C/O)$ for the high-pressure gas. We also show the rough range of abundance ratios from Pascale et al. (2023) as gray shaded regions.
4.1 Fiducial results
We first go through our “fiducial” simulations (§ 3.4).
Fig. 5 visualizes the simulation M3e7_R70_Z0.2. The first row shows the gas surface density and the distribution of VMSs at several times elapsed in simulation. At $\sim 0.59\,\rm Myr$, the GMC fragments into dense clumps, causing early star formation (green dots) at off-center locations. At $\sim 2\,\rm Myr$, the cloud further collapses and several stellar sub-clusters hierarchically form. At $\sim 3.7\,\rm Myr$, more sub-clusters form and merge at the center of the cloud complex. At this stage, strong stellar feedback drives the surrounding gas envelope outward, expanding the cloud even further at $\sim 4.3\,\rm Myr$ in the rightmost panel.
The second row shows the surface density of enriched (current minus initial) nitrogen. Recall that our simulations start with equal-mass gas cells, each with mass $m_{j}$ ($j$ is the indexing of gas cells) and metallicity $X_{i,j}$ for element $i$ (homogeneous for our case here); then after star formation, the mass of element $i$ in the same gas cell may be enriched by stellar chemical feedback, which is now $m_{j}^{\prime}X_{i,j}^{\prime}$. We then define the “enriched element mass” as $\Delta m_{i,j}=m_{j}^{\prime}X_{i,j}^{\prime}-m_{j}X_{i,j}$, and evaluate the corresponding surface density $\Delta\Sigma_{\rm N}$. With ongoing star formation, stars return nitrogen to the cloud: at $\sim 2\,\rm Myr$, hot spots of $\Delta\Sigma_{\rm N}$ correspond to the star-forming regions in the top panel (in contrast to the nearly blank map at $\sim 0.59\,\rm Myr$). Such chemical feedback is even stronger at $\sim 3.7\,\rm Myr$, leading to the density peak in $\Delta\Sigma_{\rm N}$ at the center of the cluster. However, the enriched gas has been pushed outward by stellar feedback at $\sim 4.3\,\rm Myr$.
@@ 4.1 Fiducial results
The last row shows the abundance ratios N/O, C/O and He/H. They do not deviate significantly from the initial GMC values except for a late time at $\sim 4.3\,\rm Myr$. This enrichment occurs rapidly in just $\lesssim 0.6\,\rm Myr$ given the null signals at $\sim 3.7\,\rm Myr$. In particular, $\log(\rm N/O)$ increases from $\sim-1.2$ up to $\sim 0$, similar to the level of enrichment observed for the high-pressure gas of the Sunburst LyC cluster (Pascale et al., 2023).
Interestingly, although the second row of Fig. 5 shows that nitrogen expelled by the stars is widespread in the cloud complex, the last row indicates that strongly elevated N/O (as well as elevated C/O and He/H) shows within $\sim 10\,$pc of the super star cluster. Significantly high N/O values only arise in a fraction of the nitrogen-enriched gas, which will be studied later in this section.
@@
4.1.1 Super star cluster formation
The left two columns of Fig. 6 characterize the evolution of the simulated star cluster. For each simulation, we identify the onset time of star formation $t_{0}$, which is used to define the cluster age $t-t_{0}$. The exact time $t_{0}$ is different in each simulation but is typically around $\sim 0.1\,\rm Myr$, suggesting that star formation is triggered soon after GMC collapse. The mass of newly-formed stars initially increases as $M_{\star}\propto(t-t_{0})^{2}$ (Li et al., 2018), but the growth becomes steeper later after $t-t_{0}\sim 0.3\,\rm Myr$. Moreover, star formation culminates at $\sim 1\,\rm Myr$, when half of the stellar mass ($\sim 5\times 10^{6}\,M_{\odot}$) forms. Star formation is completed at $\sim 2-3\,\rm Myr$, roughly within two global free-fall times of the natal cloud.
@@ 4.1.1 Super star cluster formation
We also plot the time evolution of the total mass of VMSs in the same panels. The total mass of VMSs first grows with time with a slope similar to the total stellar mass (but is smaller by a factor of $\sim 0.1$ as expected from the IMF), and then reaches the peak of $\sim 5\times 10^{5}\,M_{\odot}$ at $\sim 3\,$Myr, before it declines as a result of VMS mass loss and death. At $t-t_{0}\sim 4\,\rm Myr$, the VMS total mass declines to $\sim 3\times 10^{5}\,\rm M_{\odot}$, which implies $\sim 2000$ VMSs by assuming the IMF-averaged VMS mass of $\sim 150\,M_{\odot}$. Meštrić et al. (2023) find that the presence of $\sim 400$ VMSs is needed to explain the VMS stellar wind features in the Sunburst LyC cluster, significantly less than the cumulative number of VMSs expected from the default IMF and formed in our simulation. This may or may not reflect a genuine effect on the IMF, and we caution about the uncertainty in predicting the number of surviving VMSs present in the LyC cluster, due to the uncertainty in VMS evolution and cluster age.
The second column of Fig. 6 shows the time evolution of the UV half-light radius for all the stars in the field. This radius is evaluated with respect to the center of the most massive central cluster (where luminosity also peaks), following its definition by counting the UV luminosity of each star particle (but radiation transfer by the ISM is ignored). To estimate the UV luminosity of a single star, we first obtain the bolometric luminosity $L_{\rm bol}(m)$ from the luminosity-mass relation in Hopkins et al. (2018a), then the radius $R_{\star}(m)$ from the radius-mass relation in Tout et al. (1996), which further implies the effective surface temperature $T_{\rm eff}\approx 5770\,{\rm K}\,[(L_{\rm bol}/R_{\star}^{2})/(L_{\odot}/R_{\odot}^{2})]^{1/4}$. The UV luminosity of the star is then $L_{\rm UV}=f_{\rm UV}L_{\rm bol}$, where $f_{\rm UV}$ is the energy fraction of the UV band ($3.444<h\nu/{\rm eV}<8$, following FIRE definitions) in a black-body radiation spectrum $B_{\nu}(T_{\rm eff})$. We note that UV luminosity is largely dominated by massive stars, e.g., $L_{\rm UV}/m\approx 0.08\,L_{\odot}/M_{\odot}$ when $m=1\,M_{\odot}$, while it monotonically rises to $L_{\rm UV}/m\approx 7250\,L_{\odot}/M_{\odot}$ when $m=300\,M_{\odot}$. Moreover, we show the half-mass radius of the most massive cluster.
@@
These results suggest that the Sunburst LyC cluster may consist of many low-mass satellite sub-clusters surrounding a central massive compact cluster. This could explain the resolved outskirts of UV emission (Pascale et al., 2023; Vanzella et al., 2022). In our simulations, the satellites originate from the same natal cloud and are coeval with the central cluster.
Fig. 7 also shows the spatial distribution of VMSs, which are concentrated at the center of each sub-cluster, which is expected since dynamical friction is faster for the more massive stars (Chandrasekhar, 1943). Such an effect is modeled in our simulation with a tree-based algorithm (Ma et al., 2023), and aligns with the idea of mass segregation put forth in Meštrić et al. (2023). Moreover, the oblate shape of a subset of sub-clusters in Fig. 7 may suggest significant spins of young star clusters formed in turbulent GMCs.
@@
4.1.2 Chemical enrichment
The right two columns of Fig. 6 show N/O and C/O ratios in the high-pressure gas, which is selected for gas cells with $p>10^{9}\rm K\,\rm cm^{-3}$ (following Pascale et al., 2023), and the curves in these panels represent the median $\log(\rm N/O)$ and $\log(\rm C/O)$ values of these gas cells. Compared with the observed abundance ratios for the Sunburst LyC cluster from Pascale et al. (2023) (gray shaded regions), we find that the fiducial simulations can reproduce the observed N/O and C/O abundance ratios during the period $2\,{\rm Myr}<t-t_{0}\lesssim 4\,\rm Myr$. It is also possible that $\log(\rm N/O)$ can even rise above 0, before the curves break or disappear at $\sim 4\,\rm Myr$ (which suggests absence of high-pressure gas). As also indicated in Fig. 5, during this time, stellar feedback becomes effective and continuously pushes the gas envelope outward, eventually leading to depletion of the high-pressure gas.

### v2

@@ III.4 Simulation suite
The star formation physics in this study follows FIRE-3 (HopkinsWetzelWheeler_2023MNRAS.519.3154H), which evaluates the GMC star formation rate while modeling the unresolved IMF with the hybrid method described in § III.1. A gas cell forms a FIRE SSP particle only if (1) its density exceeds $1000\,\rm cm^{-3}$, (2) it is self-gravitating against turbulence, thermal, and magnetic support, (3) it exhibits convergent inflow ($\nabla\cdot\mathbf{v}<0$), and (4) its Jeans mass is smaller than the cell mass (GrudicGuszejnovHopkins_2021MNRAS.506.2199G). The star formation rate is further weighted by the cell’s molecular fraction (HopkinsWetzelKeres_2018MNRAS.480..800H). Once an SSP forms, Poisson sampling determines the number of VMSs, which are spawned as individual particles (see Shi_inprep).
@@
Radiative (UV, optical, IR) and kinetic (winds and SNe) feedback follow HopkinsWetzelKeres_2018MNRAS.477.1578H; HopkinsWetzelKeres_2018MNRAS.480..800H. Radiation transfer is treated as in HopkinsQuataertMurray_2012MNRAS.421.3488H, where the luminosity $L_{\nu}$ in each band is locally attenuated by gas and dust before propagating into the ISM, producing momentum coupling and radiative heating/cooling. In the UV, where massive stars dominate, the flux-weighted opacity is $\kappa_{\nu}=1800\,(Z/Z_{\odot})\,{\rm cm}^{2}\,{\rm g}^{-1}$, assuming dust abundance scales with metallicity. VMS luminosities in each band are computed assuming blackbody spectra, following GrudicGuszejnovHopkins_2021MNRAS.506.2199G.
Metals are treated as passive scalars, evolved via the advection–diffusion equation with stellar yield source terms, following HopkinsWetzelKeres_2018MNRAS.480..800H and RennehanBabulHopkins_2019MNRAS.483.3810R, and verified by ColbrookMaHopkins_2017MNRAS.467.2421C. In addition to resolved mixing through the continuity equation, unresolved small-scale transport is modeled with the Smagorinsky turbulent diffusion prescription (Smagorinsky_1963MWRv...91...99S; EscalaWetzelKirby_2018MNRAS.474.2194E).
The initial conditions follow previous setups (GrudicHopkinsFaucher-Giguere_2018MNRAS.475.3511G; Shi_inprep): a spherical, uniform, non-rotating GMC with turbulent velocities containing 50% solenoidal modes and total kinetic energy equal to the gravitational binding energy, $\alpha_{\rm vir}\!\equiv\!2T/|U|\!=\!2$. Although observations show a broad $\alpha_{\rm vir}$ distribution (e.g., SunLeroySchinnerer_2020ApJ...901L...8S), this value is typical for massive GMCs (ChevanceKrumholzMcLeod_2023ASPC..534....1C). The initial magnetic field is uniform, with magnetic energy set to 1% of the gravitational binding energy. The initially uniform density and field quickly become turbulent owing to the imposed velocity field. Elemental abundances follow the empirical relations in Eqs. (3) and (4).
@@
Each simulation is labeled M%g_R%g_Z%g, denoting the GMC mass $M_{\rm cloud}$ ($M_{\odot}$), radius $R_{\rm cloud}$ (pc), and metallicity $Z$ ($Z_{\odot}$); e.g., M3e7_R70_Z0.2 corresponds to $M_{\rm cloud}=3\times 10^{7}\,M_{\odot}$, $R_{\rm cloud}=70\,{\rm pc}$, and $Z=0.2\,Z_{\odot}$. All GMCs start with $64^{3}$ equal-mass gas cells and evolve for 10 Myr. When VMSs are included, we adopt a Kroupa IMF extending to $m_{\rm max}=300\,M_{\odot}$, with stars above $m_{\rm cut}=75\,M_{\odot}$ treated individually and less massive ones modeled as FIRE SSPs. We run four simulation groups for different goals (Table 1).
@@
    Fiducial simulations. We use the FIRE+VMS module explained earlier to perform the fiducial simulations. Based on properties of the LyC leaking cluster (PascaleDaiMcKee_2023ApJ...957...77P), our fiducial simulations have fixed $M_{\rm cloud}=3\times 10^{7}\,M_{\odot}$ but vary $R_{\rm cloud}$ ($60,70,80\,\rm pc$) and $Z/Z_{\odot}$ (0.2, 0.15, 0.25). As will be explained in § IV, these simulations can (roughly) reproduce the observations.
@@
    GMC surface density tests. This variation keeps the fiducial setup but initializes with different $M_{\rm cloud}\,(10^{7},10^{8}\,\rm M_{\odot})$, while keeping the same $R_{\rm cloud}$ ($60,70,80\,\rm pc$) and $Z/Z_{\odot}=0.2$. These simulations explore the impact of the GMC initial mass (and surface density), to accommodate the order-of-unity uncertainty in the star cluster’s mass due to the magnification factor of gravitational lensing (PascaleDaiMcKee_2023ApJ...957...77P).
@@
    FIRE-only simulations. These simulations remove explicitly-modeled VMSs and utilize FIRE-3 physics (where the underlying IMF only extends to $100\,M_{\odot}$) to highlight the effect of VMSs in shaping the star formation, especially the chemical evolution. They initialize with $M_{\rm cloud}=3\times 10^{7}\,M_{\odot}$, $R_{\rm cloud}=60,70,80\,\rm pc$, and $Z=0.2\,Z_{\odot}$. In particular, since we removed “failed SNe” from FIRE-3 when implementing the hybrid FIRE+VMS model (§ III.3), we also run such a version of simulations in contrast to the standard FIRE-3 physics.
@@
IV Results
Figure 7: Star formation and the chemical abundance of high-pressure ($10^{9}\,\rm K\,cm^{-3}$) gas throughout the evolution of our “fiducial” group of simulations, as a function of “time since star formation.” These simulations initialize with different GMCs (M3e7_R60, M3e7_R70, M3e7_R80) and metallicities: $0.2\,Z_{\odot}$ (top row), $0.15\,Z_{\odot}$ (second row), and $0.25\,Z_{\odot}$ (bottom row). Each row, from left to right, shows the 1) star formation history, in general (solid), or of the most massive bound cluster (dotted), or of all VMSs (dot dashed); 2) UV half-light radius (of all stars in the field; solid) and the 3D half-mass radius (of the most massive bound cluster; dotted); 3) median $\log(\rm N/O)$ for the high-pressure gas; 4) median $\log(\rm C/O)$ for the high-pressure gas. We also show the rough range of abundance ratios from PascaleDaiMcKee_2023ApJ...957...77P as gray shaded regions.
IV.1 Fiducial results
We first go through our “fiducial” simulations (§ III.4).
Figure 6 visualizes the M3e7_R70_Z0.2 simulation. The top row shows the gas surface density and VMS distribution at several times. By $\sim 0.59\,\rm Myr$, the GMC fragments into dense clumps, forming stars (green dots) off-center. At $\sim 2\,\rm Myr$, hierarchical sub-clusters develop during further collapse, and by $\sim 3.7\,\rm Myr$, they merge near the cloud center. Strong stellar feedback then expels the surrounding gas, driving large-scale expansion by $\sim 4.3\,\rm Myr$.
The second row shows the surface density of enriched nitrogen, $\Delta\Sigma_{\rm N}$. Each gas cell of mass $m_{j}$ and initial abundance $X_{i,j}$ (for element $i$) evolves to $m_{j}^{\prime}$ and $X_{i,j}^{\prime}$ after star formation and feedback. The enriched mass is defined as $\Delta m_{i,j}=m_{j}^{\prime}X_{i,j}^{\prime}-m_{j}X_{i,j}$, from which we compute $\Delta\Sigma_{\rm N}$. As stars form, nitrogen enrichment becomes visible: at $\sim 2\,\rm Myr$, $\Delta\Sigma_{\rm N}$ peaks at star-forming regions (contrast the nearly blank map at $\sim 0.6\,\rm Myr$); by $\sim 3.7\,\rm Myr$, feedback enhances central enrichment, and by $\sim 4.3\,\rm Myr$, the enriched gas is driven outward.
@@ IV.1 Fiducial results
The bottom row shows the abundance ratios N/O, C/O, and He/H. They remain close to the initial GMC values except at late times ($\sim 4.3\,\rm Myr$), when enrichment occurs rapidly within $\lesssim 0.6\,\rm Myr$, given the absence of signals at $\sim 3.7\,\rm Myr$. Notably, $\log(\mathrm{N/O})$ rises from $\sim-1.2$ to $\sim 0$, consistent with the enrichment level observed in the high-pressure gas of the Sunburst LyC cluster (PascaleDaiMcKee_2023ApJ...957...77P).
Although the second row of Fig. 6 shows nitrogen widely dispersed throughout the cloud, the bottom row indicates that strongly elevated N/O (and similarly enhanced C/O and He/H) appears only within $\sim 10$ pc of the super star cluster. Such extreme N/O values arise in only a small fraction of the enriched gas, a point explored further below.
@@
IV.1.1 Super star cluster formation
The left two columns of Fig. 7 characterize the evolution of the simulated star cluster. For each simulation, we identify the onset time of star formation $t_{0}$, which is used to define the cluster age $t-t_{0}$. The exact time $t_{0}$ is different in each simulation but is typically around $\sim 0.1\,\rm Myr$, suggesting that star formation is triggered soon after GMC collapse. The mass of newly-formed stars initially increases as $M_{\star}\propto(t-t_{0})^{2}$ (LiKleinMcKee_2018MNRAS.473.4220L), but the growth becomes steeper later after $t-t_{0}\sim 0.3\,\rm Myr$. Moreover, star formation culminates at $\sim 1\,\rm Myr$, when half of the stellar mass ($\sim 5\times 10^{6}\,M_{\odot}$) forms. Star formation is completed at $\sim 2-3\,\rm Myr$, roughly within two global free-fall times of the natal cloud.
@@ IV.1.1 Super star cluster formation
We also plot the time evolution of the total mass of VMSs in the same panels. The total mass of VMSs first grows with time with a slope similar to the total stellar mass (but is smaller by a factor of $\sim 0.1$ as expected from the IMF), and then reaches the peak of $\sim 5\times 10^{5}\,M_{\odot}$ at $\sim 3\,$Myr, before it declines as a result of VMS mass loss and death. At $t-t_{0}\sim 4\,\rm Myr$, the VMS total mass declines to $\sim 3\times 10^{5}\,\rm M_{\odot}$, which implies $\sim 2000$ VMSs by assuming the IMF-averaged VMS mass of $\sim 150\,M_{\odot}$. MestricVanzellaUpadhyaya_2023A&A...673A..50M find that the presence of $\sim 400$ VMSs is needed to explain the VMS stellar wind features in the Sunburst LyC cluster, significantly less than the cumulative number of VMSs expected from the default IMF and formed in our simulation. This may or may not reflect a genuine effect on the IMF, and we caution about the uncertainty in predicting the number of surviving VMSs present in the LyC cluster, due to the uncertainty in VMS evolution and cluster age.
The second column of Fig. 7 shows the time evolution of the UV half-light radius for all the stars in the field. The radius is measured relative to the center of the most massive cluster (which also marks the luminosity peak), defined by summing the UV luminosities of all star particles while neglecting ISM radiative transfer. For each star, the bolometric luminosity $L_{\rm bol}(m)$ follows the luminosity–mass relation of HopkinsWetzelKeres_2018MNRAS.480..800H, and the stellar radius $R_{\star}(m)$ follows ToutPolsEggleton_1996MNRAS.281..257T. The effective temperature is $T_{\rm eff}\approx 5770\,{\rm K}\,[(L_{\rm bol}/R_{\star}^{2})/(L_{\odot}/R_{\odot}^{2})]^{1/4}$. The UV luminosity is $L_{\rm UV}=f_{\rm UV}L_{\rm bol}$, where $f_{\rm UV}$ is the fraction of black-body emission in the FIRE UV band ($3.444<h\nu/{\rm eV}<8$). Massive stars dominate the UV output: $L_{\rm UV}/m\approx 0.08\,L_{\odot}/M_{\odot}$ at $1\,M_{\odot}$, increasing to $L_{\rm UV}/m\approx 7250\,L_{\odot}/M_{\odot}$ at $300\,M_{\odot}$. We also show the half-mass radius of the most massive cluster.
@@
These results suggest that the Sunburst LyC cluster may consist of many low-mass satellite sub-clusters surrounding a central massive compact cluster. This could explain the resolved outskirts of UV emission (PascaleDaiMcKee_2023ApJ...957...77P; VanzellaCastellanoBergamini_2022A&A...659A...2V). In our simulations, the satellites originate from the same natal cloud and are coeval with the central cluster.
Fig. 8 also shows the spatial distribution of VMSs, which are concentrated at the center of each sub-cluster, which is expected since dynamical friction is faster for the more massive stars (Chandrasekhar_1943ApJ....97..255C). Such an effect is modeled in our simulation with a tree-based algorithm (MaHopkinsKelley_2023MNRAS.519.5543M), and aligns with the idea of mass segregation put forth in MestricVanzellaUpadhyaya_2023A&A...673A..50M. Moreover, the oblate shape of a subset of sub-clusters in Fig. 8 may suggest significant spins of young star clusters formed in turbulent GMCs.
@@
IV.1.2 Chemical enrichment
The right two columns of Fig. 7 show N/O and C/O ratios in the high-pressure gas, which is selected for gas cells with $p>10^{9}\rm K\,\rm cm^{-3}$ (following PascaleDaiMcKee_2023ApJ...957...77P), and the curves in these panels represent the median $\log(\rm N/O)$ and $\log(\rm C/O)$ values of these gas cells. Compared with the observed abundance ratios for the Sunburst LyC cluster from PascaleDaiMcKee_2023ApJ...957...77P (gray shaded regions), we find that the fiducial simulations can reproduce the observed N/O and C/O abundance ratios during the period $2\,{\rm Myr}<t-t_{0}\lesssim 4\,\rm Myr$. It is also possible that $\log(\rm N/O)$ can even rise above 0, before the curves break or disappear at $\sim 4\,\rm Myr$ (which suggests absence of high-pressure gas). As also indicated in Fig. 6, during this time, stellar feedback becomes effective and continuously pushes the gas envelope outward, eventually leading to depletion of the high-pressure gas.